# Risk Limited Asset Allocation with a Budget Threshold Utility Function and Leptokurtotic Distributions of Returns

### v1

@@ 5. Maximum Expected Utility Subject to an Exogenous Risk-Limit
6. Interpretation of the Result
I believe that despite the apparent naïvety of the problem proposed the solution is, in fact, quite interesting. It is well known that a risk-limited gross-profit maximizing investor should obtain a position as large as possible in the direction of the alpha[giller2023essays]. The introduction of the budget threshold, as specified in this work, changes such a long-short, or “binary,” trading algorithm to a three state system that considers long, short, and flat positions only. The term $\sigma\tau(\kappa)$ can be thought of as a “risk cost” that must be exceeded to make risk-taking worthwhile to the investor.
@@ 6. Interpretation of the Result
(11) $s^{2}=\mathbb{V}[r]=\sigma^{2}\times\frac{2^{2\kappa}\Gamma(3\kappa)}{\Gamma(\kappa)}.$
@@
(12) $\sigma\tau(\kappa)=\frac{1}{4}\sqrt{\frac{\Gamma(\kappa)}{\Gamma(3\kappa)}}\frac{\Gamma(2\kappa+1}{\Gamma(\kappa+1)}\times s.$
@@
Nevertheless, given the commonly acknowledged non-stationarity of financial distributions and the simplicity of the approach, it seems well motivated to follow a “semi-empirical” approach, similar to that adopted in Nuclear Physics[bethe1936semf, weizsacker1935semf], and replace this parametric expression with a simple constant, $K$, to be determined from empirical optimization (i.e. backtesting). A practical algorithm for traders is then
@@
where $\alpha$ and $s$ are the mean and standard deviation of future returns under the conditional distribution used by a trader for a given asset.

### v2

@@ 5. Maximum Expected Utility Subject to an Exogenous Risk-Limit
6. Interpretation of the Result and
a Practical Algorithm for Traders
I believe that, despite the apparent naïvety of the problem proposed, the solution is, in fact, quite interesting. It is well known that a risk-limited gross-profit maximizing investor should obtain a position as large as possible in the direction of the alpha[giller2023essays]. The introduction of the budget threshold, as specified in this work, changes such a long-short, or “binary,” trading algorithm to a three state system that considers long, short, and flat positions only. The term $\sigma\tau(\kappa)$ can be thought of as a “risk cost” that must be exceeded to make risk-taking worthwhile to the investor.
@@ 6. Interpretation of the Result and
(11) $s^{2}=\mathbb{V}[r]=\frac{2^{2\kappa}\Gamma(3\kappa)}{\Gamma(\kappa)}\times\sigma^{2}.$
@@
(12) $\sigma\tau(\kappa)=\frac{1}{4}\sqrt{\frac{\Gamma(\kappa)}{\Gamma(3\kappa)}}\frac{\Gamma(2\kappa+1)}{\Gamma(\kappa+1)}\times s.$
@@
Nevertheless, given the commonly acknowledged non-stationarity of financial distributions and the simplicity of the optimal algorithm, it seems well motivated to follow a “semi-empirical” approach, similar to that adopted in Nuclear Physics[bethe1936semf, weizsacker1935semf], and replace this parametric expression with a simple constant, $K$, to be determined from empirical optimization (i.e. backtesting). A practical algorithm for traders is then
@@
where $\alpha$ and $s$ are the mean and standard deviation of future returns under the conditional distribution used by a trader for a given asset, and $K\approx 0.4$. This trading algorithm, in the form of a “holding function,” [^8] Literally a function that tells the trader want position to hold in response to their inputs. is illustrated in figure 4.
Figure 4. The “holding function” corresponding to the optimal strategy for a trader with a budget threshold utility function expressed as the relative position size as a function of the standardized alpha.